OpenStudy (anonymous):

log base 2(x-1)+log base 2(x+1)=3, what is x and why. I get two different answers.

OpenStudy (anonymous):

write out that left side in terms of one log: logbase2 [(x-1)(x+1)] = 3 so now you can write this out in exponential form: 2^3 = (x-1)(x+1)

OpenStudy (anonymous):

solve this quadratic..

OpenStudy (anonymous):

this how you got your answer? and what are your answers?

OpenStudy (anonymous):

one answer was 2 and the other answer was 3. I get when using your method

OpenStudy (anonymous):

I get 3 using your method

OpenStudy (anonymous):

well, solving that quadratic yields two solutions but 1 of those solutions need to be thrown out because it is not a solution to the original equation. so there is only 1 solution.. x=3.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

yw... :)